dukedadon:

solve this equation. 2(x+4)-3x=7-5x

dukedadon:

@vocaloid help please?

TETSXPREME:

@extrinix

axie:

Ok lets get started, the steps are 1. Distribute 2.Combine like terms 3. Rearange terms 4. Solve for X

axie:

Intrinix is going to help with those steps!

Extrinix:

Ok so, first off you need to distribute \(2(x+4)-3x=7-5x\) \(2x+8-3x=7-5x\) Combine like terms \(-1x+8=7-5x\) Add 5x to both sides \(4x+8=7\) Subtract 8 from both sides \(4x=-1\) Divide by 4 on both sides \(x=-\dfrac{1}{4}\)

dukedadon:

@extrinix wrote:
Ok so, first off you need to distribute ty \(2(x+4)-3x=7-5x\) \(2x+8-3x=7-5x\) Combine like terms \(-1x+8=7-5x\) Add 5x to both sides \(4x+8=7\) Subtract 8 from both sides \(4x=-1\) Divide by 4 on both sides \(x=-\dfrac{1}{4}\)